OpenStudy (anonymous):

Damien wants to deposit $1200 into savings accounts at three different banks: Bank of Chance, Merchant Bank, and Utopian Financing. He will deposit three times as much into Merchant Bank as Bank of Chance because they offer a higher interest rate. He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits. How much will Damien deposit into the Utopian Financing savings account? $240 $250 $200 $750 Please help. If you explain how you got the answer i will give you a medal

OpenStudy (anonymous):

No. I'm not sure how to do that

OpenStudy (anonymous):

@johnweldon1993 can you help??

OpenStudy (johnweldon1993):

Hey sorry I wasn't on yesterday...but in case you still need the help Think of it like this...assign every bank a variable Let Bank of Chance = C Merchant Bank = M Utopian Financing = U We know he is going to make deposit 1200 total into these 3 banks...so \[\large C + M + U = 1200\] "He will deposit three times as much into Merchant Bank as Bank of Chance " This means \[\large 3M = C\] there will be 3 times more than C "He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits" So the total of the other 2 deposits are C + M And f U will be 20% of that...we have \[\large .20(C+M)\] So now, we have 3 equations \[\large C + M + U = 1200\] \[\large 3M = C\] \[\large U = .20(C + M)\] Focus on the last 2 equations first... We know that C = 3M so plug that into the last equation \[\large U = .20(3M + M)\] \[\large U = .20(4M)\] Now notice....the first equation....we have \[\large C + M + U = 1200\] Well we now know that U = .20(4M) and that C = 3M lets plug all that in \[\large 3M + M + (.20(4M)) = 1200\] \[\large 4M + (.20(4M) = 1200\] \[\large 4M + \frac{4}{5}M = 1200\] \[\large \frac{24}{5}M = 1200\] \[\large M = 250\] He will deposit 250 into M Now that you know that...plug in that into the 2nd equation \[\large 3M = C\] \[\large 3(250) = C\] \[\large 750 = C\] So 750 into chance.. and finally for U...we know that \[\large C+M+U = 1200\] \[\large 750 + 250 + U = 1200\] \[\large 1000 + U = 1200\] \[\large U = 200\] He will deposit 200 into Utopian

OpenStudy (anonymous):

Thank you!! i'm still working on this module so this helps a lot. and i have a different scenario about taxes and things. could you maybe help with that too? @johnweldon1993